### v1

@@ Abstract
One-shot voice conversion (VC) aims to alter the timbre of speech from a source speaker to match that of a target speaker using just a single reference speech from the target, while preserving the semantic content of the original source speech. Despite advancements in one-shot VC, its effectiveness decreases in real-world scenarios where reference speeches, often sourced from the internet, contain various disturbances like background noise. To address this issue, we introduce Noro, a Noise Robust One-shot VC system. Noro features innovative components tailored for VC using noisy reference speeches, including a dual-branch reference encoding module and a noise-agnostic contrastive speaker loss. Experimental results demonstrate that Noro outperforms our baseline system in both clean and noisy scenarios, highlighting its efficacy for real-world applications. Additionally, we investigate the hidden speaker representation capabilities of our baseline system by repurposing its reference encoder as a speaker encoder. The results shows that it is competitive with several advanced self-supervised learning models for speaker representation under the SUPERB settings, highlighting the potential for advancing speaker representation learning through one-shot VC task.
{IEEEkeywords}
Noise-robust voice conversion, one-shot voice conversion, speaker representation learning.
@@ 1 INTRODUCTION
One-shot voice conversion (VC) changes the timbre of speech from a source speaker to that of a target speaker using just one reference speech sample from the target, while maintaining the original speech’s semantic content. This task has been extensively researched in the deep learning era, employing various methods and producing promising results [1, 2, 3, 4, 5, 6]. However, the strong performance of these one-shot VC methods is mainly demonstrated in controlled academic settings, where high-quality data is used for both training and evaluation. As noted in [7, 8, 9, 10], their effectiveness significantly decreases in more challenging real-world scenarios, where reference speeches often collected from the internet are characterized by various interferences, such as background noise, which leads to a drop in the quality and similarity of the converted speech.
@@
Few previous works tried to address this challenge to achieve noise-robust VC. Some researchers [11, 12] attempt to integrate a pre-trained or jointly trained speech enhancement (SE) module into standard VC systems to improve robustness. However, these methods are not end-to-end and inevitably increase computational costs during inference. Other researchers [8, 13, 10] suggest strategies during training phase, such as direct data augmentation [8], to mitigate the impact of noise. Despite these efforts, there is still room for improvement, especially in extremely noisy conditions where the Signal-to-Noise Ratio (SNR) is below 5 dB [13].
@@
To address these challenges, we introduce a novel Noise-Robust One-shot VC system named Noro. Noro is built on a one-shot VC baseline system based on diffusion [14], which generates speech from pitch and semantic representations obtained through a source encoder and speaker timbre representations via a reference encoder. To achieve noise-robustness of the system, we focus on learning speaker timbre representations that are agnostic to noise. We hypothesize that maintaining noise-agnostic input representations for the acoustic model will prevent its performance from deteriorating due to noise interference.
@@
Therefore, we propose the following designs: First, we introduce a novel dual-branch reference encoding module. This module consists of a dual-branch transformer encoder structure with shared weights: one branch encodes clean reference speech, while the other encodes its noisy counterpart generated through data augmentation. Next, we devise a noise-agnostic contrastive speaker loss to maximize the similarity between samples (whether clean or noisy) from the same speaker while minimizing it for those from different speakers. This loss ensures that the dual-branch reference encoding module learns to represent speaker timbre independent of the acoustic environment of the reference speeches. Experimental results demonstrate that Noro significantly improves the robustness of our baseline system in both clean and noisy environments, making it an ideal choice for real-world applications that encounter noisy reference speeches.
@@
Our contributions in this work can be summarized as follows.
  • We proposed Noro, a Noise-Robust One-shot VC system. It achieves noise robustness through innovative designs, including a dual-branch reference encoding module, a noise-agnostic contrastive speaker loss. Experimental results demonstrate that Noro significantly outperforms state-of-the-art one-shot VC systems in noisy scenarios, establishing it as the preferred choice for real-world applications requiring reliable, high-quality VC in diverse and challenging acoustic environments.
  • We investigated the hidden speaker representation capabilities of the proposed one-shot VC system by repurposing its reference encoder as a speaker encoder (VC-SPK2VEC). VC-SPK2VEC exhibits comparable performance with several SOTA SSL models in speaker representation under the SUPERB setting, highlighting the overlooked potential for speaker representation learning via the one-shot VC task.
The paper is organized as follows: Section 2 discusses the related works. Section 3 introduces the methodology, including a detailed description of our baseline one-shot voice conversion (VC) system and the proposed Noro system. Section 4 first presents the experimental setup and results, evaluating the performance of Noro under various acoustic conditions and comparing it to baseline systems. We then explore the impact of different source encoder models on the performance of Noro. Subsequently, we explore the potential of re-purposing the reference encoder from our VC system as a self-supervised speaker encoder. Finally, Section 5 provides concluding remarks, summarizing the findings and highlighting future research directions.
2 Related work
2.1 Noise-robust Voice Conversion
2.1.1 One-shot Voice Conversion
Numerous studies have sought high-quality one-shot voice conversion using methods such as auto-encoders [22, 23, 24, 25, 26, 1, 27, 28, 29], normalizing flows [30], and self-supervised learning (SSL) models [31, 32, 2, 33, 34]. For example, Diff-VC [1] uses an encoder that predicts an “average voice,” transforming mel features of each phoneme into averaged mel features across a large multi-speaker dataset, followed by decoding the mel for converted speech. Similarly, FreeVC [28] performs voice conversion by extracting disentangled content and target speaker information, then reconstructs the waveform using an end-to-end VITS framework. Most recently, FaCodec-VC [29] employs a factorized codec approach to achieve state-of-the-art one-shot voice conversion performance. It effectively disentangles speech attributes into content, prosody, and timbre representations using an information bottleneck, enabling strong one-shot voice conversion by altering the timbre of the source speech to match the target speaker. Despite these advances, the effectiveness of such models decreases in challenging real-world scenarios, where reference speech samples from the internet often contain interferences like background noise, reducing the quality and similarity of the converted speech.
To address the issue of noise robustness, several studies have been conducted. For instance, Valentini-Botinhao et al.[11] proposed an RNN-based speech enhancement method to improve noise robustness, utilizing LSTM networks for two-stage processing of input signals to effectively suppress noise. Similarly, Chan et al.[12] developed a speech enhancement (SE)-assisted VC system, where the VC and SE components are jointly optimized to produce high-quality converted speech signals. Additionally, Miao et al. [35] introduced a noise-robust voice conversion method that enhances high-quefrency components through sub-band cepstrum conversion and fusion. Their approach leverages bidirectional long short-term memory networks to convert vocal tract parameters from a source to a target speaker in noisy conditions. However, these approaches are not end-to-end, requiring separate stages for enhancement and synthesis, which adds complexity and increases computational overhead during inference. To address this, some researchers have proposed training strategies to improve robustness. Huang et al.[8] proposed applying random degraded data augmentations during training to force the model to filter out degradations and perform VC accurately. Xue et al.[10] introduced a noise-controllable WaveGAN that learns noise-independent acoustic representations to improve conversion quality in noisy environments. Du et al. [13] employed domain adversarial training to ensure robustness against various noise conditions during conversion. Despite these advancements, significant challenges remain, especially in extremely noisy environments where the Signal-to-Noise Ratio (SNR) is below 5 dB. This paper proposes the Noro system to address this challenge, introducing a dual-branch reference encoding module and a noise-agnostic contrastive speaker loss, which significantly enhance noise robustness even in challenging acoustic environments.
2.2 Self-supervised Learning for Speaker Representation Learning
Various self-supervised learning (SSL) models have been proposed for speech processing tasks. These models are generally categorized into three types: generative, discriminative, and multi-task models [17]. Generative models, such as APC [36], aim to reconstruct masked or future frames, focusing on predicting the missing or upcoming parts of the speech sequence. Discriminative models like wav2vec [18] and HuBERT [20] use contrastive learning to distinguish between different segments of speech. Multi-task models, such as WavLM+ [21], incorporate multiple pretraining objectives to learn versatile representations for downstream tasks. These models have shown their effectiveness in various aspects of speech processing, including content, speaker, semantics, and paralinguistic information [17]. Inspired by [15, 16], this study further explores the speaker representation capabilities of one-shot VC systems. The motivation arises from the observation that the reference encoder within one-shot VC systems, designed to capture the vocal timbre of diverse speakers—including previously unseen ones—naturally functions as a speaker encoder. This aligns well with self-supervised speaker representation learning.
@@ 3 Methodology
In this section, we first introduce our baseline one-shot VC system. As depicted in Fig. 1 (a), this system comprises a source encoder to simultaneously encode semantic and pitch representations (Sec. 3.1.1), a reference encoder to extract speaker timbre representation from the reference speech (Sec. 3.1.2), and a diffusion model that utilizes these representations as conditions to predict the mel-spectrogram of the target speech (Sec. 3.1.3). To achieve noise-robustness against noisy reference speeches, we introduce Noro. Fig. 1 (b) provides an overview of Noro. Noro builds upon the baseline system, but replaces its original reference encoders with a dual-branch reference encoding module and a noise-agnostic contrastive speaker loss.
Figure 1: Overview of the model architecture of our baseline system and Noro.
3.1 Our Baseline System
@@ 3.1.1 Source Encoder
To derive semantic representations from the source speech, our baseline system uses a pre-trained and frozen HuBERT model [20] as the semantic extractor. This model encodes the source speech into continuous embeddings. We then apply K-Means quantization to these embeddings. Instead of using K-Means cluster IDs (discrete semantic tokens), we use the quantized continuous embeddings, which preserve more information, for semantic representation. The resulting semantic representation for the source speech is denoted as $s_{src}$.
For pitch representations, we use the open-source software PyWorld [^1] https://github.com/JeremyCCHsu/Python-Wrapper-for-World-Vocoder to extract frame-level F0 values from the source speech. We normalize these values and use them for pitch representation, denoted as $p_{src}$.
After extracting the semantic and pitch representations, we employ a Conformer [37] encoder to simultaneously encode $s_{src}$ and $p_{src}$, obtaining the final source representation $h_{src}$.
@@ 3.1.2 Reference Encoder
To encode the speaker timbre in the reference speech, we use a cross-attention scheme similar to the speech prompting mechanism in NaturalSpeech 2 [38]. Specifically, we first employ a Transformer based reference encoder to convert the melspectrogram of the reference speech into a hidden sequence. Instead of allowing the diffusion model to directly attend to this hidden sequence, we employ two attention blocks.
In the first attention block, we use $m$ randomly initialized query embeddings to attend to the hidden sequence, resulting in $h_{ref}$, a hidden sequence of length $m$ that represents the reference speaker’s timbre at the utterance level. In the second attention block, the output hidden sequence of the WaveNet Layer serves as the query, while $h_{ref}$ functions as both the key and value. The attention results of the second attention block are then used as conditional information for a FiLM layer in the final acoustic modeling process.
@@ 3.1.3 Diffusion Model
To enhance in-context learning for improved one-shot generation, each training sample is randomly segmented during the training phase. A segment comprising 25-45% of the total sample is designated as the reference speech, while the remaining portion functions as both the source and target speech. We utilize a WaveNet-based diffusion model [38], denoted as $s_{\theta}$, to predict the melspectrogram $z_{tgt}$ of the target speech.
We describe both the diffusion (forward) process and the denoising (reverse) process using the stochastic differential equation. In the forward process, Gaussian noise is added to the melspectrogram $z_{tgt}$:
[…]
where $\beta_{t}$ represents the noise schedule function, $\epsilon$ denotes randomly sampled Gaussian noise, and $t\in[0,1]$. The reverse process can be formulated as:
$dz_{tgt}^{t}=-\left(\frac{1}{2}z_{tgt}^{t}+\nabla\log p_{t}(z_{tgt}^{t})\right)\beta_{t}\,dt+\sqrt{\beta_{t}}\epsilon.$ (2)
$s_{\theta}$ takes the noised melspectrogram $z_{tgt}^{t}$, the time step $t$, and the conditions of the source representation $h_{src}$, containing both semantic and pitch information, and the reference representation $h_{ref}$, containing speaker timbre information, to estimate $\nabla\log p_{t}(z_{tgt}^{t})$. The loss function of the diffusion model can be written as:
$L_{diff}=\left\|s_{\theta}(z_{tgt}^{t},t,h_{src},h_{ref})-\nabla\log p_{t}(z_{tgt}^{t})\right\|_{1}.$ (3)
During the inference stage, we gradually denoise $z_{tgt}^{t}$ utilizing the estimated score $s_{\theta}(z_{tgt}^{t},t,s_{src},p_{src},h_{ref})$. This process initiates with Gaussian noise, initially sampled as $z_{tgt}^{1}$.
3.2 Noro
@@ 3.2.1 Dual-branch Reference Encoding Module
To improve the noise robustness of the baseline system, we replaced its original reference encoder with a dual-branch reference encoding module. Unlike the baseline system, which uses a single reference encoder, this module incorporates two reference encoders sharing identical model weights to form a dual-branch structure.
For each clean training sample, we randomly mix clean reference speech with eight types of noise from the DEMAND database [39] at SNRs following a normal distribution (0,20) dB to create noisy reference speeches. Subsequently, the clean and noisy reference speeches are processed by separate reference encoders using the same query embeddings. This process generates clean and noisy reference representations $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$, respectively.
Considering the weight-sharing scheme used in this dual-branch reference encoding module, the method described here qualifies as a ”training strategy” for enhancing our baseline system. Initially, we load the Noro system with weights from a pre-trained baseline system. During the training phase, the average of $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$ is fed into the diffusion model for acoustic modeling and acts as the keys and values in the second attention block. During the inference phase, only one reference encoder is utilized, and the Noro system’s structure mirrors that of the baseline system.
@@ 3.2.2 Noise-agnostic Contrastive Speaker Loss
After obtaining the representations of the clean speeches $h_{\text{ref}}$ and their noisy counterparts $h^{\prime}_{\text{ref}}$, we introduce a noise-agnostic contrastive speaker loss. This loss function aims to maximize the similarity between samples (clean or noisy) from the same speaker while minimizing it for samples from different speakers.
First, as $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$ are hidden sequences of length $m$, we perform average pooling over the length dimension to form a comprehensive reference representation that captures the reference speaker’s timbre at the utterance level. We then concatenate $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$ along the batch size dimension:
$h_{\text{all}}=\left[h_{\text{ref}};h^{\prime}_{\text{ref}}\right].$ (4)
Given that the noisy speeches share the same speakers as the clean ones, we can express $y_{\text{all}}$ as:
$y_{\text{all}}=\left[y_{\text{spk}};y_{\text{spk}}\right].$ (5)
Finally, our proposed noise-agnostic contrastive speaker loss is formulated as:
$\mathcal{L}_{\text{ref}}=\frac{1}{2N}\sum_{i=1}^{2N}\text{CrossEntropy}\left(\frac{h_{i}\cdot h_{j}^{T}}{\tau},M_{i,j}\right),$ (6)
where $h_{i}$ and $h_{j}$ are the $i$-th and $j$-th reference representations in $h_{\text{all}}$, $\tau$ is the temperature parameter adjusting the scaling of the logits, $N$ is the batch size, and the mask matrix $M_{i,j}$ is defined as follows:
$M_{i,j}=\begin{cases}1&\text{if }y_{i}=y_{j},\\ 0&\text{otherwise},\end{cases}$ (7)
where $y_{i}$ and $y_{j}$ are the speaker labels for $h_{i}$ and $h_{j}$ respectively. This loss ensures that the reference encoders represent the vocal timbre of different speakers regardless of noise interference, thus enhancing the robustness of the system.
The total loss function for the Noro system is defined as follows:
$L_{total}=\alpha L_{diff}+\beta L_{ref},$ (8)
where $\alpha$ and $\beta$ are the weights assigned to each loss component.

### v2

@@ Abstract
The effectiveness of one-shot voice conversion (VC) decreases in real-world scenarios where reference speeches, which are often sourced from the internet, contain various disturbances like background noise. To address this issue, we introduce Noro, a noise-robust one-shot VC system. Noro features innovative components tailored for VC using noisy reference speeches, including a dual-branch reference encoding module and a noise-agnostic contrastive speaker loss. Experimental results demonstrate that Noro outperforms our baseline system in both clean and noisy scenarios, highlighting its efficacy for real-world applications. Additionally, we investigate the hidden speaker representation capabilities of our baseline system by repurposing its reference encoder as a speaker encoder. The results show that it is competitive with several advanced self-supervised learning models for speaker representation under the SUPERB settings, highlighting the potential for advancing speaker representation learning through one-shot VC tasks.
11footnotetext: Equal contribution. Listed in alphabetical order.
@@ 1 Introduction
One-shot voice conversion (VC) alters the timbre of speech from a source speaker to that of a target speaker using just one reference speech sample from the target speaker, while maintaining the semantic content of the original speech. This task has been extensively studied in the deep learning era, employing various methods and producing promising results [1, 2, 3, 4, 5]. However, the strong performance of these one-shot VC methods is mainly demonstrated in controlled academic settings, where high-quality data is used for both training and evaluation. As noted in [6, 7, 8, 9], their effectiveness significantly decreases in more challenging real-world scenarios, where reference speeches often collected from the internet are characterized by various interferences, such as background noise, leading to suboptimal converted speech.
@@
Few works have addressed this challenge to achieve noise-robust VC. Some researchers [10, 11] attempt to integrate a pre-trained or jointly trained speech enhancement module into standard VC systems to improve robustness. However, these methods are not end-to-end and inevitably increase computational costs during inference. Other researchers [7, 12, 9, 13] suggest training strategies, such as data augmentation [7], to mitigate the impact of noise. However, there is still room for improvement, especially in extremely noisy conditions where the Signal-to-Noise Ratio (SNR) is below 5 dB [12].
@@
To address these challenges, we introduce a novel noise-robust one-shot VC system named Noro. Noro is built on a one-shot VC baseline system based on diffusion. This model is trained to generate speech from pitch and semantic representations obtained through a source encoder and speaker timbre representations via a reference encoder. To achieve noise-robustness with noisy reference speeches, we focus on learning speaker timbre representations that are agnostic to noise. We hypothesize that maintaining noise-agnostic input representations for the acoustic model will prevent its performance from deteriorating due to noise interference.
@@
Therefore, we propose the following designs: Firstly, we introduce a dual-branch reference encoding module. This module consists of two transformer encoders with shared weights: one branch encodes clean reference speech, while the other encodes its noisy counterpart generated through data augmentation. Next, we devise a noise-agnostic contrastive speaker loss to maximize the similarity between samples (whether clean or noisy) from the same speaker while minimizing it for those from different speakers. This loss ensures that the dual-branch reference encoding module learns to represent speaker timbre independent of the acoustic environment of the reference speeches. Experimental results show that Noro significantly enhances the robustness of our baseline system in diverse noisy environments, with only a minor reduction in performance in clean settings, making it well-suited for real-world applications involving noisy reference speeches.
@@ 2 Methodology
In this section, we first introduce our baseline one-shot VC system and Noro. As depicted in Fig. 1, Noro builds upon the baseline system, but replaces its original reference encoders with a dual-branch reference encoding module and a noise-agnostic contrastive speaker loss.
Figure 1: Model architecture of our baseline system and Noro.
2.1 The Design of Our Baseline System
@@ 2.1.1 Source Encoder
To derive semantic representations from the source speech, our baseline system employs a pre-trained and frozen HuBERT model [19, 21] as the semantic extractor. This model encodes the source speech into continuous embeddings. We then apply K-Means quantization to these embeddings. Instead of using K-Means cluster IDs (discrete semantic tokens), we use the quantized continuous embeddings, which preserve more information, for semantic representation. The resulting semantic representation for the source speech is denoted as $s_{\text{src}}$. For pitch representations, we use the open-source software PyWorld to extract frame-level F0 values from the source speech. We apply min-max normalization to these values and use them for pitch representation, denoted as $p_{\text{src}}$. After extracting the semantic and pitch representations, we employ a convolutional neural network (CNN)-based encoder to simultaneously encode $s_{\text{src}}$ and $p_{\text{src}}$, obtaining the final source representation $h_{\text{src}}$.
@@ 2.1.2 Reference Encoder
To encode the speaker timbre in the reference speech, we use a cross-attention scheme similar to the speech prompting mechanism in NaturalSpeech 2 [22]. Specifically, we first employ a Transformer-based reference encoder to convert the melspectrogram of the reference speech into a hidden sequence. Instead of allowing the diffusion model to directly attend to this hidden sequence, we employ two attention blocks. In the first attention block, we use $m$ randomly initialized query embeddings to attend to the hidden sequence, resulting in $h_{\text{ref}}$, a hidden sequence of length $m$ that represents the reference speaker’s timbre at the utterance level. In the second attention block, the output hidden sequence of the WaveNet Layer serves as the query, while $h_{\text{ref}}$ functions as both the key and value. The attention results of the second attention block are then used as conditional information for a FiLM layer in the final acoustic modeling process.
@@ 2.1.3 Diffusion Model
To enhance in-context learning for improved one-shot generation, each training sample is randomly segmented during the training phase. A segment comprising 25-45% of the total sample is designated as the reference speech, while the remaining portion functions as both the source and target speech. We utilize a WaveNet-based diffusion model [22], denoted as $s_{\theta}$, to predict the melspectrogram $\mathbf{z}_{\text{tgt}}$ of the target speech. We describe both the diffusion (forward) process and the denoising (reverse) process using the stochastic differential equation. In the forward process, Gaussian noise is added to the melspectrogram $\mathbf{z}_{\text{tgt}}$: $\mathbf{z}_{\text{tgt}}^{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta_{s}ds}\mathbf{z}_{\text{tgt}}+(\mathbf{I}-e^{-\int_{0}^{t}\beta_{s}ds})\boldsymbol{\epsilon},$ where $\beta_{t}$ represents the noise schedule function, $\boldsymbol{\epsilon}$ denotes randomly sampled Gaussian noise, and $t\in[0,1]$. The reverse process can be formulated as: $d\mathbf{z}_{\text{tgt}}^{t}=-\left(\frac{1}{2}\mathbf{z}_{\text{tgt}}^{t}+\nabla\log p_{t}(\mathbf{z}_{\text{tgt}}^{t})\right)\beta_{t}\,dt+\sqrt{\beta_{t}}\boldsymbol{\epsilon}.$
$s_{\theta}$ takes the noised melspectrogram $\mathbf{z}_{\text{tgt}}^{t}$, the time step $t$, and the conditions of the source representation $h_{\text{src}}$, containing both semantic and pitch information, and the reference representation $h_{\text{ref}}$, containing speaker timbre information, to estimate $\nabla\log p_{t}(\mathbf{z}_{\text{tgt}}^{t})$. The loss function of the diffusion model can be written as: $\mathcal{L}_{\text{diff}}=\left\|s_{\theta}(\mathbf{z}_{\text{tgt}}^{t},t,h_{\text{src}},h_{\text{ref}})-\nabla\log p_{t}(\mathbf{z}_{\text{tgt}}^{t})\right\|_{1}.$ During the inference stage, we gradually denoise $\mathbf{z}_{\text{tgt}}^{t}$ utilizing the estimated score $s_{\theta}(\mathbf{z}_{\text{tgt}}^{t},t,s_{\text{src}},p_{\text{src}},h_{\text{ref}})$. This process initiates with Gaussian noise, initially sampled as $\mathbf{z}_{\text{tgt}}^{1}$.
2.2 The Design of Noro
@@ 2.2.1 Dual-branch Reference Encoding Module
To improve the noise robustness of the baseline system, we replaced its original reference encoder with a dual-branch reference encoding module. This module consists of two reference encoders sharing identical model weights as a dual-branch structure. For each clean training sample, we randomly mix clean reference speech with eight types of noise from the DEMAND database [23] at SNRs following a normal distribution (0,20) dB to create noisy reference speeches. Subsequently, the clean and noisy reference speeches are processed by separate reference encoders using the same query embeddings. This process generates clean and noisy reference representations $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$, respectively. Considering the weight-sharing scheme used in this dual-branch reference encoding module, the method described here qualifies as a “training strategy” for enhancing our baseline system. Initially, we load the Noro system with weights from a pre-trained baseline system. During the training phase, the average of $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$ is fed into the diffusion model for acoustic modeling and acts as the keys and values in the second attention block. During the inference phase, only one reference encoder is utilized, and the Noro system’s structure mirrors that of the baseline system.
@@ 2.2.2 Noise-agnostic Contrastive Speaker Loss
After obtaining the representations of the clean speeches $h_{\text{ref}}$ and their noisy counterparts $h^{\prime}_{\text{ref}}$, we introduce a noise-agnostic contrastive speaker loss. This loss function aims to maximize the similarity between samples (clean or noisy) from the same speaker while minimizing it for samples from different speakers. First, as $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$ are hidden sequences of length $m$, we perform average pooling over the length dimension to form a comprehensive reference representation that captures the reference speaker’s timbre at the utterance level. We then concatenate $h_{\text{ref}}$ and $h^{\prime}_{\text{ref}}$ along the batch size dimension: $h_{\text{all}}=\left[h_{\text{ref}};h^{\prime}_{\text{ref}}\right].$ Given that the noisy speeches share the same speakers as the clean ones, $y_{\text{all}}=\left[y_{\text{spk}};y_{\text{spk}}\right].$ Finally, our proposed noise-agnostic contrastive speaker loss is formulated as: $\mathcal{L}_{\text{ref}}=\frac{1}{2N}\sum_{i=1}^{2N}\text{CrossEntropy}\left(\frac{h_{i}\cdot h_{j}^{T}}{\tau},\mathbf{M}_{i,j}\right),$ where $h_{i}$ and $h_{j}$ are the $i$-th and $j$-th reference representations in $h_{\text{all}}$, $\tau$ is the temperature parameter adjusting the scaling of the logits, $N$ is the batch size, and the mask matrix $\mathbf{M}_{i,j}$ is defined as follows: $\mathbf{M}_{i,j}=\begin{cases}1&\text{if }y_{i}=y_{j},\\ 0&\text{otherwise},\end{cases}$ where $y_{i}$ and $y_{j}$ are the speaker labels for $h_{i}$ and $h_{j}$ respectively. This loss ensures that the reference encoders represent the vocal timbre of different speakers regardless of noise interference, thus enhancing the robustness of the system. The total loss function for the Noro system is defined as follows: $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{diff}}+\beta\mathcal{L}_{\text{ref}}$, where $\alpha$ and $\beta$ are the weights assigned to each loss component.